This question refers to: Laws of motion
A mass of 2 kg travelling at 2 m/s is struck by a mass of 4 kg travelling at 3 m/s. If the coefficient of restitution for the two materials is 0.4, calculate their final velocities.
Give the final velocity value in m/s of the 2kg mass here:
1
Expert's answer
2014-12-24T02:15:19-0500
Given:
m1=2
v1b=2 (velosity before collision)
m2=4
v2b=3 (velosity before collision)
cr=0.3 (coeffitient of restitution)
We want to find v1a (velosity after collision)-?
Solution:
According ofComversation of the impulse
m1*v1b - m2*v2b = - m1*v1a + m2*v2a
and
definition of coeffitient of restitution
cr= (v1b-v1a)/(v2a-v2b)
we have
v1a= [(m1*v1b-m2*v2b)-(m2*v2+m2*v1b/cr)]/[m1-m2/cr]=3.4 m/s
Answer: v1a=3.4 m/s